# Virne: A Comprehensive Benchmark for Deep RL-based Network Resource Allocation in NFV

### v1

@@ D.4 Scalability on Network Sizes
D.4.1 Performance on Large-scale Networks
@@
We observe that PPO-DualGAT and PPO-DualGCN, continue to exhibit strong performance on the large-scale WX500 network. This demonstrates the good scalability of their learned policies to more complex physical infrastructures. While other RL methods with simpler extractors exhibit limited performance, this suggests that 11 the capability to effectively capture and process complex topological information becomes increasingly critical in larger networks. Additionally, while MCTS and meta-heuristics (GA-Meta, PSO-Meta, SA-Meta, TS-Meta) can sometimes achieve competitive solution quality, their AST becomes a significant bottleneck on larger networks.
D.4.2 Analysis of Solving Time Scale
@@
Figure 9 illustrates these relationships for various algorithms. Notably, we observe that 12 rL-based methods generally maintain a much flatter and lower AST curve, while meta-heuristics exhibit a substantial increase. Their inference time is primarily dictated by the forward pass through the trained neural network, which scales less dramatically with VN size compared to exhaustive search or numerous iterations. For other heuristics, they are typically the fastest, exhibiting very low and stable ASTs irrespective of VN size.
The Virne benchmark, through systematic AST profiling, starkly highlights the trade-off between solution quality and computational scalability. While MCTS and certain meta-heuristics might achieve competitive solution quality in some scenarios (as shown in the main paper), their substantial computational overhead, especially with increasing problem dimensions, limits their applicability in dynamic, large-scale NFV environments. 13 RL-based methods, particularly those utilizing GNNs, strike a more favorable balance. Simpler RL methods and traditional heuristics offer the highest computational speed but may compromise on optimality.
D.5 Validation on Emerging Network
@@
D.5.1 Heterogeneous Resourcing Networks
@@
As illustrated in Table 9, we observe that PPO-DualGAT and PPO-GAT&S2S demonstrate superior performance in heterogeneous resourcing environments, outperforming PPO-DualGCN. 14 The attention mechanism and edge awareness in GATs, which can weigh the importance of different features and neighbors, appear particularly adept at navigating multi-dimensional resource constraints. Simpler RL architectures like PPO-MLP and PPO-GCN, while competent, tend to show a slight dip in performance compared to scenarios with homogeneous resources. 15 Effectively managing and balancing multiple, diverse resource requirements simultaneously poses a greater challenge for architectures that may not explicitly differentiate or dynamically weigh these varied resource dimensions.
D.5.2 Latency-aware Edge Networks
Edge computing networks bring computational resources closer to end-users, making them ideal for latency-sensitive applications. In such scenarios, NFV-RA algorithms must not only allocate resources efficiently but also strictly adhere to the delay requirements of services. This evaluation uses the WX100 topology with a VN arrival rate $\eta=0.08$, which could represent a typical edge deployment with specific traffic characteristics. The critical factor introduced here is a latency constraint for each virtual link. When a virtual link is mapped to a physical path, the cumulative propagation delay of the physical links constituting that path must not exceed the virtual link’s specified maximum tolerable latency, i.e., 100ms. This necessitates that the algorithms should possess the ability to manage constraint management. We additionally consider the latency attributes as the input features of neural networks.
As illustrated in Table 9, the introduction of latency constraints generally leads to lower RAC and LRC values across all algorithms compared to scenarios without such strict path requirements. However, PPO-DualGAT again shows the most robust performance, achieving the highest RAC (68.20%), LRC (0.714), and LAR (12056.58). In contrast, 16 many traditional heuristics and MCTS find it more challenging to satisfy both resource and latency constraints simultaneously, often resulting in more pronounced performance degradation.
@@ Appendix E Discussion on Future Directions
The empirical analysis enabled by the Virne benchmark illuminates several promising avenues for advancing deep RL-based NFV-RA methods. While current leading algorithms demonstrate significant capabilities, addressing the following challenges could lead to even more robust, efficient, and practical solutions.
Representation Learning for Cross-graph Status
@@
Future research should focus on developing more sophisticated representation learning techniques to capture the intricate, dynamic interplay between VN requirements and PN states, including the evolving mapping status itself. While Virne shows the strength of dual-GNN architectures in processing VN and PN features separately then combining them, there’s a need for models that can learn even richer cross-graph relational embeddings and path-level attribute awareness [72, 73]. This includes better capturing dependencies across multiple resource types and complex constraints, potentially leading to more nuanced and context-aware embedding decisions. Techniques that can explicitly model the partially embedded state of VNs and the resultant constraints on future placements are crucial.
Robust Learning for Complex and Hard Constraints
@@
The management of multifaceted and often conflicting constraints (e.g., resource capacity, latency, reliability, energy consumption) remains a central challenge in NFV-RA, which requires guaranteeing zero-violation. Thus, it is significant to explore constraint-aware RL frameworks that more explicitly model and navigate these hard constraints. This could involve advancing constrained policy optimization methods and developing novel reward structures that better reflect constraint satisfaction [74, 50]. The goal is to train safe policies that are not only effective in optimizing primary objectives but are also inherently robust in satisfying diverse operational constraints.
Scalability in Extremely Large-scale Networks
@@
Virne’s scalability studies demonstrate that while current RL methods, particularly GNN-based ones, scale better than many traditional approaches, their computational and memory demands can still grow significantly with network size. Addressing NFV-RA in truly massive, carrier-grade networks requires further breakthroughs in algorithmic scalability. Future directions could include exploring hierarchical RL where policies operate at different levels of abstraction, building a non-autoregressive solution modeling method [60], or designing GNN architectures that are less sensitive to the global network scale [75]. Methods that can learn transferable knowledge or localizable policies that effectively stitch together solutions for very large infrastructures are highly desirable [18].
Generalizable Policy Cross Varying Scale
@@
Achieving policies that generalize effectively not only to unseen network topologies but also across varying scales (both PN and VN sizes) and highly dynamic operational conditions (e.g., non-stationary demand patterns, unexpected resource outages) is a critical frontier. Virne’s generalization experiments provide a baseline, but future research should investigate advanced techniques such as meta-RL [76] to train agents that can quickly adapt to new VNR sizes or types, curriculum learning to incrementally expose agents to more complex scenarios, and domain randomization to enhance robustness against a wider array of unseen conditions [62, 19]. The ultimate aim is to develop omni-generalizable NFV-RA solutions that require minimal retraining for deployment in diverse and evolving network environments.

### v2

@@ D.3 Scalability on Network Sizes
D.3.1 Performance on Large-scale Networks
@@
We observe that PPO-DualGAT and PPO-DualGCN continue to exhibit strong performance on the large-scale WX500 network. This demonstrates the good scalability of their learned policies to more complex physical infrastructures. While other RL methods with simpler extractors, like PPO-MLP, exhibit limited performance, this suggests that 14 the capability to effectively capture and process complex topological information becomes increasingly critical in larger networks. Additionally, while MCTS and meta-heuristics (GA-Meta, PSO-Meta, SA-Meta, TS-Meta) can sometimes achieve competitive solution quality, their AST becomes a significant bottleneck on larger networks. Furthermore, it is interesting that 15 there is also a trade-off between maximizing immediate revenue and ensuring long-term network efficiency. Simpler models like PPO-MLP exemplify the first approach, achieving the highest total revenue (LAR) by aiming to maximize acceptance volume. In contrast, sophisticated models like PPO-DualGAT exemplify the second, securing the highest resource efficiency (LRC) with a strategic policy that prioritizes high-quality placements to preserve resource utilization efficiency.
D.3.2 Analysis of Solving Time Scale
@@
Figure 12 illustrates these relationships for various algorithms. We observe that rRL-based methods generally maintain a much flatter and lower AST curve, while meta-heuristics exhibit a substantial increase. Their inference time is primarily dictated by the forward pass through the trained neural network, which scales less dramatically with VN size compared to exhaustive search or numerous iterations. For other heuristics, they are typically the fastest, exhibiting very low and stable ASTs irrespective of VN size. While MCTS and certain meta-heuristics might achieve competitive solution quality in some scenarios, their substantial computational overhead, especially with increasing problem dimensions, limits their applicability in dynamic, large-scale NFV environments. This further highlights that 16 as the problem space grows exponentially, explicit search strategies at decision time become computationally prohibitive, rendering them infeasible for dynamic environments that require near-instantaneous decisions. Additionally, 17 RL-based methods, particularly those utilizing GNNs, strike a more favorable balance. Simpler RL methods and traditional heuristics offer the highest computational speed but may compromise on optimality.
Overall, through systematic AST profiling analysis, Virne starkly highlights the trade-off between solution quality and computational scalability.
D.4 Validation on Emerging Network
@@
D.4.1 Heterogeneous Resourcing Networks
@@
As illustrated in Table 9, we observe that PPO-DualGAT and PPO-GAT&S2S demonstrate superior performance in heterogeneous resourcing environments, outperforming PPO-DualGCN. In heterogeneous resource scenarios, GAT-based models (e.g., PPO-DualGAT and PPO-GAT&S2S) excel because they can dynamically weigh the importance of different resource types, such as CPU versus a scarcer GPU, which is a more nuanced approach than the uniform feature treatment in GCNs. This highlights that 18 attention mechanisms over features and neighbors are highly effective for multi-dimensional node-level constraints. Simpler RL architectures like PPO-MLP and PPO-GCN, while competent, tend to show a slight dip in performance compared to scenarios with homogeneous resources. This reveals that 19 Effectively managing and balancing multiple, diverse resource requirements simultaneously poses a greater challenge for architectures that may not explicitly differentiate or dynamically weigh these varied resource dimensions.
D.4.2 Latency-aware Edge Networks
Edge computing networks bring computational resources closer to end-users, making them ideal for latency-sensitive applications. In such scenarios, NFV-RA algorithms must not only allocate resources efficiently but also strictly adhere to the delay requirements of services. This evaluation uses the WX100 topology with a VN arrival rate $\eta=0.08$, which could represent a typical edge deployment with specific traffic characteristics. The critical factor introduced here is a latency constraint for each virtual link. When a virtual link is mapped to a physical path, the cumulative propagation delay of the physical links constituting that path must not exceed the virtual link’s specified maximum tolerable latency, i.e., 100ms. This necessitates that the algorithms should possess the ability to manage constraint management. We additionally consider the latency attributes as the input features of neural networks for all deep RL-based methods.
As illustrated in Table 9, the introduction of latency constraints generally leads to lower RAC and LRC values across all algorithms compared to scenarios without such strict path requirements. However, PPO-DualGAT again shows the most robust performance, achieving the highest RAC (68.20%), LRC (0.714), and LAR (12056.58). In contrast, traditional heuristics and MCTS struggle to satisfy resource and latency simultaneously, suffering larger drops. This reveals that 20 The fundamental weakness of traditional heuristics is exposed in scenarios with path-level constraints. Their decoupled "rank nodes, then find paths" approach is brittle, as an optimal node choice based on local metrics can make it impossible to subsequently find a valid low-latency path, leading to cascading failures.
@@ Appendix E Discussion on Future Directions
The empirical analysis enabled by the Virne benchmark illuminates several promising avenues for advancing deep RL-based NFV-RA methods. Although state-of-the-art algorithms show substantial capability, significant challenges remain that hinder more robust, efficient, and practical solutions. In this section, we outline several key future directions, and we provide an empirical study of an emerging algorithm addressing some of these directions in Figure 13.
Figure 13: Performance comparison of PPO-DualGAT and its advanced variations that are augmented by the safe RL method proposed in Wang et al. (2025b) and the meta-RL proposed in Wang et al. (2024c) training methods. Please see their MDP formulation in Appendix C.2. The results show that both the safe RL and meta-RL variants consistently outperform the basic algorithm, with the safe RL approach achieving higher improvements. This underscores the importance of robust constraint management and generalization in solving complex resource allocation problems, which are highlighted in this section for future algorithmic innovation.
E.1 Representation Learning for Cross-graph Status
@@
Future research should focus on developing more sophisticated representation learning techniques to capture the intricate, dynamic interplay between VN requirements and PN states, including the evolving mapping status itself. While Virne demonstrates the strength of dual-GNN architectures in processing VN and PN features separately, there is significant potential in exploring emerging architectures such as Transformers and diffusion models. These advanced models could enable the learning of even richer cross-graph relational embeddings and path-level attribute awareness (Wang et al., 2019; Guo et al., 2025). Specifically, they may offer superior capabilities in capturing long-range dependencies across multiple resource types and handling complex constraints, potentially leading to more nuanced and context-aware embedding decisions. Techniques that explicitly model the partially embedded state of VNs and the resultant constraints on future placements are crucial, and Virne’s extensible architecture provides a ready platform for benchmarking these novel designs.
E.2 Robust Learning for Complex and Hard Constraints
@@
The management of multifaceted and often conflicting constraints (e.g., resource capacity, latency, reliability, energy consumption) remains a central challenge in NFV-RA, which requires guaranteeing zero-violation. Thus, it is significant to explore constraint-aware RL frameworks that more explicitly model and navigate these hard constraints. This could involve advancing constrained policy optimization methods and developing novel reward structures that better reflect constraint satisfaction (Gu et al., 2022; Wang et al., 2025b). The goal is to train safe policies that are not only effective in optimizing primary objectives but are also inherently robust in satisfying diverse operational constraints.
E.3 Scalability in Extremely Large-scale Networks
@@
Virne’s scalability studies demonstrate that while current RL methods, particularly GNN-based ones, scale better than many traditional approaches, their computational and memory demands can still grow significantly with network size. Addressing NFV-RA in truly massive, carrier-grade networks requires further breakthroughs in algorithmic scalability. Future directions could include exploring hierarchical RL where policies operate at different levels of abstraction, building a non-autoregressive solution modeling method (Bengio et al., 2021), or designing GNN architectures that are less sensitive to the global network scale (Wang et al., 2024a). Methods that can learn transferable knowledge or localizable policies that effectively stitch together solutions for very large infrastructures are highly desirable (Geng et al., 2023).
E.4 Generalizable Policy Cross Varying Scenarios
@@
Achieving policies that generalize effectively not only to unseen network topologies but also across varying scales (both PN and VN sizes) and highly dynamic operational conditions (e.g., non-stationary demand patterns, unexpected resource outages) is a critical frontier. To facilitate research in this direction, Virne explicitly incorporates a Multi-Task MDP framework (see Appendix C.2.1), allowing users to define diverse tasks encompassing different physical topologies, resource distributions, and traffic patterns. As demonstrated by the meta-learning results in Figure 13, this foundation enables the training of agents capable of adapting to new environments. Future research should leverage this flexible task definition to further investigate advanced techniques such as meta-RL (Beck et al., 2023), curriculum learning to incrementally expose agents to more complex scenarios, and domain randomization to enhance robustness against a wider array of unseen conditions (Zhou et al., 2023; Wang et al., 2024c). The ultimate aim is to develop omni-generalizable NFV-RA solutions that require minimal retraining for deployment in diverse and evolving network environments.
Appendix F Limitations on Sim-to-Real Gap
As a simulation-based benchmark, Virne makes specific abstractions to ensure the computational efficiency required for training Deep RL agents. There are several primary gaps between our default simulation and physical deployment:
  • Simulation Granularity (Flow vs. Packet). Virne functions as a flow-level orchestration simulator. To maintain high training throughput, it abstracts packet-level dynamics (e.g., microsecond queuing delays, jitter) into statistical link attributes. While this simplifies latency modeling, it allows for rapid algorithmic iteration.
  • Infrastructure Observability (Global vs. Partial). The default experimental settings assume a centralized SDN controller with global state visibility. In contrast, real-world large-scale networks may only offer partial observability. Virne mitigates this via a modular architecture, allowing users to apply feature masking to simulate POMDP scenarios without altering the core engine.
  • Resource Dynamics (Idealized vs. Noisy). The default environment assumes deterministic resource allocation. Physical systems, however, often experience fluctuations due to background processes or hardware overhead ("noisy neighbors"). Virne supports extensibility in its transition function, allowing users to inject stochastic noise to model these environmental perturbations.
Appendix G The Use of Large Language Models Statement
The authors use Large Language Models (LLMs) as an assistive tool in the preparation of this manuscript, in accordance with the ICLR 2026 policy. The core roles of LLMs are described as follows. First, we utilize the LLM-powered search and summarization tools to survey related works and understand the current state of the field. Second, we use LLM-based coding assistants to assist in coding and suggest optimizations, thereby accelerating the development of the Virne codebase. Third, while preparing the manuscript, we use LLMs to proofread, check grammar, and refine the language in the manuscript for improved clarity and readability.